I need help to understand and explain to an 11 year old how to solve 32y + 10 − 2 = 24
Answer by lynnlo(4176)

PROBLEM:===32Y+10-2=24
(1)ADD FRACTIONS OR INTEGERS TOGETHER
32Y+8=24
(2)MOVE "-24" TO THE LEFT
32Y+8-24=0
(3)ADD FRACTIONS OR INTEGERS TOGETHER
32Y+ -16=0
REMOVE SIGN IN FRONT OF "-16"
32Y-16=0
SOLVE LINEAR EQUATION
32Y-16=0
EQUIVALENT TO 32*Y-16=0
SOLUTION
Y=0,5 OR Y=1/2